A two digit number has its digits sum equal to 11.
If you reverse the digits of this number, the new two digit number is 27 more than the original two digit number.
I know that the original number is 47 and the new digits in reverse is 74 (which is 27 more than the original 47)
My question is a solution, including all algebraic steps to solve this problem.
Thanks
ML

Answer by scott8148(6628)

let x="tens digit" and y="units digit" __ x+y=11 __ x=11-y

10y+x=10x+y+27

substituting __ 10y+(11-y)=10(11-y)+y+27 __ 9y+11=137-9y __ 18y=126